Define the relative refractive index difference for an optical fiber and show how it

May be related to the numerical aperture. A step index fiber with a large core diameter compared with the wavelength of the transmitted light has an acceptance angle in air of 22° and a relative refractive index difference of 3%. Estimate the numerical aperture and the critical angle at the core–cladding interface for the fiber.



1
Expert's answer
2022-03-15T16:49:42-0400

"\Delta =1-\frac{n_2}{n_1},"

"n_2=0.97n_1,"

"n_2=1,"

"n_1=1.03,"

"NA=\sqrt{n_1^2-n_2^2}=0.25,"

"\beta=\arcsin\frac {n_1}{\sin\alpha}=25°."